OpenStudy (anonymous):

(x+3)^1/2-1=x Please help me solve that

OpenStudy (anonymous):

you need help still

OpenStudy (anonymous):

Find the partial-fraction decomposition of the following expression: (x^2 + 1) / [ x (x - 1)^3 ] The factor x – 1 occurs three times in the denominator. I will account for that by forming fractions containing increasing powers of this factor in the denominator, like this: (x^2 + 1)/[x(x - 1)^3] = A/(x - 1) + B/(x - 1)^2 + C/(x - 1)^3 + D/x Now I multiply through by the common denominator to get: x2 + 1 = Ax(x – 1)2 + Bx(x – 1) + Cx + D(x – 1)3 I could use a system of equations to solve for A, B, C, and D, but the other method seemed easier. The two zeroing numbers are x = 1 and x = 0: so x = 1: 1 + 1 = 0 + 0 + C + 0, so C = 2 x = 0: 1 = 0 + 0 + 0 – D, so D = –1 But what do I do now? I have two other variables, namely A and B, for which I need values. But since I've got values for C and D, I can pick any two other x-values, plug them in, and get a system of equations that I can solve for A and B. The particular x-values I choose aren't important, so I'll pick smallish ones: x = 2: Copyright © Elizabeth Stapel 2006-2011 All Rights Reserved (2)2 + 1 = A(2)(2 – 1)2 + B(2)(2 – 1) + (2)(2) + (–1)(2 – 1)3 4 + 1 = 2A + 2B + 4 – 1 5 = 2A + 2B + 3 1 = A + B x = –1: (–1)2 + 1 = A(–1)(–1 – 1)2 + B(–1)(–1 – 1) + (2)(–1) + (–1)(–1 – 1)3 1 + 1 = –4A + 2B – 2 + 8 2 = –4A + 2B + 6 2A – B = 2 I'm still stuck solving a system of equations, but by using the easier method to solve for C and D, I now have a simpler system to solve. Adding the two equations, I get 3A = 3, so A = 1. Then B = 0 (so that term in the expansion "vanishes"), and the complete decomposition is: 1/(x - 1) + 2/(x - 1)^3 - 1/x

OpenStudy (cwrw238):

yes - as simple as that

OpenStudy (anonymous):

Oh man…. that's confusing.. sorry @ayla2001 but would you be able to just work with me step by step?

OpenStudy (anonymous):

x=1 x in (-oo:+oo) ((x+3)^1)/2-1 = x // - x ((x+3)^1)/2-x-1 = 0 (x+3)/2-x-1 = 0 (x+3)/2+(2*(-x))/2+(-1*2)/2 = 0 x+2*(-x)-1*2+3 = 0 3-x-2 = 0 1-x = 0 (1-x)/2 = 0 (1-x)/2 = 0 // * 2 1-x = 0 1-x = 0 // - 1 -x = -1 // * -1 x = 1

OpenStudy (cwrw238):

(x + 3)^1/2 - 1 = x add 1 to both sides (x + 3)^1/2 = x + 1 square both sides x + 3 = (x + 1)^2 x + 3 = x^2 + 2x + 1 x^2 + x - 2 = 0 now solve this for x

OpenStudy (anonymous):

do you get it know

OpenStudy (anonymous):

Got it thanks. Would yo be able to help with another problem?

OpenStudy (cwrw238):

the way you wrote the problem is ambiguous - it can be interpreted in different ways